a stone is dropped from a height of 45m. what will be the distance traveeled by it during the last one second of its motion??

Grade:11

2 Answers

Hiba Musthafa
6 years ago
total time taken=(2h/g)1/2 = (2*45/10) ½  = (9) ½ = 3
distance travelled in last sec = S= u + a/2 (2n-1)
S= 0 + 10/2 (2*3 – 1)
= 25 m
 
Rawnak
4 years ago
Time taken = s=ut + 1\2at×t
  45 = 1/2×10t×t = ✓9=3
 S= u+a
( n-1/2)
0+10 ( 3 -1/2) 
25m
Use formula of nth sec